# JUREX-4E: Juridical Expert-Annotated Four-Element Knowledge Base for Legal Reasoning

### v1

@@ Appendix G A Case Study of LCR
Experts Four-elements LLM-generated Experts-guided
Four Elements
@@
Object: The ownership of property of companies, enterprises, or other organizations and the integrity of official duties. In this case, crime objetcs are the properties of the company, enterprise, or other organization where the perpetrator is employed.
Objective Aspect: The act of using one’s position for personal gain by illegally appropriating the property of the organization, where the amount involved is significant. Specifically, B used his position to embezzle X yuan from the payments of six units for personal investment and consumption.
@@
Table 10: The results of the case four elements obtained through different methods in LCR. Notes refer to expert notes related to the charge retrieved during the search. LLM-generated and Experts-guided indicate whether using JUREX-4E’s four-elements of the crime to guide LLM in generating the four elements. Red parts mean the knowledge from JUREX-4E, while blue parts show the LLM’s internal knowledge. By incorporating JUREX-4E, the model better emphasizes conviction and sentencing related information and provides more detailed descriptions of critical case facts.

### v2

@@ Appendix I A Case Study of LCR
Expert-FETs LLM-generated Expert-guided
FETs
@@
Object: The ownership of property of companies, enterprises, or other organizations and the integrity of official duties. In this case, crime objects are the properties of the company, enterprise, or other organization where the perpetrator is employed.
Objective Aspect: Using one’s position for personal gain by illegally appropriating the property of the organization, where the amount involved is significant. Specifically, B used his position to embezzle X yuan from the payments of six units for personal investment and consumption.
@@
Table 16: Results of the case-FETs using different methods for LCR. Expert-FETs refer to expert annotations of the predicted charges. LLM-generated and Expert-guided indicate whether the Expert-FETs of predicted charges are used to guide the LLM in generating the case-FETs. Red parts indicate knowledge from experts, while blue parts represent the LLM’s internal knowledge. By incorporating JUREX-4E, the model better emphasizes conviction- and sentencing-related information and provides more detailed descriptions of key case facts.